Grade 12IIT JEE Entrance Exam

Delta H(formation) of urea is given as -333.5Kj/mol
find deltaU(internal energy change) of urea

vineet chatterjee
11 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
The rformation of urea can take place as:
2 NH3(g) + CO2(g) → (NH2)2CO(s) + H2O(l)
so using the relation,
dG=dH-TdS, we would get for the chemical reactions as:
dH=dU+RTdn
where dn=number of moles in product side – number of moles in reactant side
so dH=dU-3RT
[3 comes for each one mole of urea formation, change in moles of gaseous components=0-(2+1)]
If we know the absolute temperature of this reaction occured, we can calculate the change in internal energy(dU)